Two angles are complementary. The measure of one angle is 21more than twice the measure of the other angle. Find the measures of the angles.
I am a parent trying to help a sophmore with this question. I just wanted to see if I was on the right track. Thanks for any help you could give.

Answer by rapaljer(4671)

There are (at leastl!) TWO ways to set up this equation. I think the easier way is to:

Let x= one of the angles
2x+21 = the other angle

The equation is
x + 2x +21 = 90
3x +21=90
3x = 69
x=23 degrees
2x+21 = 46+21 = 67 degrees

Check: 23 +67 = 90

The other way is to
Let x = the first angle
90-x = the second angle

The equation is from this sentence, where of course "IS" means "=":
The second angle IS 21 more than twice the first angle.
90-x= 2x+21

It comes out the same.

R^2 at SCC